OpenStudy (anonymous):

Hi my question asks Use a calculator to solve each equation on the interval 0

myininaya (myininaya):

how about you write sec(theta) in terms of cos first since the cos function is on your calc and sec isn't

OpenStudy (anonymous):

i did.

OpenStudy (anonymous):

I have to input the inverse of costheta(-4) but i get a domain error when i do it as a fraction. How do i remove fraction from 1/costheta=-4 into cosinverse theta = -4?

myininaya (myininaya):

so that means you have \[\frac{1}{\cos(\theta)}=-4 \] you need to first flip both sides then you can take arccos() of both sides

myininaya (myininaya):

are you want to cross multiply you can i will rewrite it for you in a sec \[\frac{1}{\cos(\theta)}=\frac{-4}{1}\] But I would just flip both sides (faster)

myininaya (myininaya):

flip both sides you get \[\cos(\theta)=\frac{-1}{4}\] be careful when you use arccos( ) on both sides it will only give you one number between 0 and 2pi there is actually one more x that occurs in that interval that satisfies cos(theta)=-1/4